OpenStudy (lilai3):

@RadEn please help! Write a radical expression for each square root. 4x^4y^2

OpenStudy (raden):

4x^4y^2 = {(4x^4y^2)^2}^(1/2) = sqrt(4x^4y^2)^2 = sqrt(4^2x^(4*2)y^(2*2) = sqrt(16x^8y^4)

OpenStudy (lilai3):

uhm...it looks confusing... LOL

OpenStudy (raden):

2 is squareroot of 4, right ?

OpenStudy (lilai3):

yes

OpenStudy (raden):

in othere words 2 is sqrt of 2^2 that's the principle to forms a number to radical form

OpenStudy (lilai3):